OpenStudy (anonymous):

Factor completely: 3a^2-20ab+12b^2

OpenStudy (anonymous):

\[3a ^{2}-20ab+12b ^{2}\]

jimthompson5910 (jim_thompson5910):

multiply the first and last coefficients 3*12 = 36

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 36 and that add to -20

OpenStudy (anonymous):

18 and 2

jimthompson5910 (jim_thompson5910):

actually -18 and -2 -18 times -2 = +36 -18 plus -2 = -20

OpenStudy (anonymous):

What is the next step? I'm confused on the different variables.

jimthompson5910 (jim_thompson5910):

\[\Large 3a ^{2}\color{red}{-20ab}+12b ^{2}\] \[\Large 3a ^{2}\color{red}{-18ab}\color{red}{-2ab}+12b ^{2}\] now factor by grouping

OpenStudy (anonymous):

(3a^2-18ab) (-2ab+12b^2)

jimthompson5910 (jim_thompson5910):

don't forget the plus in between (3a^2-18ab) + (-2ab+12b^2)

jimthompson5910 (jim_thompson5910):

now factor each group

OpenStudy (anonymous):

3a(a-6b) 2b(-1a+6b)

OpenStudy (anonymous):

did i do it wrong? i feel like the 2nd group is incorrect.

jimthompson5910 (jim_thompson5910):

(3a^2-18ab) + (-2ab+12b^2) 3a(a-6b) + (-2ab+12b^2) 3a(a-6b) -2b(a - 6b) there's one more step

OpenStudy (anonymous):

What would the next step be?

OpenStudy (anonymous):

oh i know it :D

OpenStudy (anonymous):

(3a-2b) (a-6b)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thanks for the help @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

no problem